OpenStudy (anonymous):

use the unit circle to find the exact value of the trig function csc 960 degrees

zepdrix (zepdrix):

Hmmmm

zepdrix (zepdrix):

So what is 960 degrees, on... earth? Like in 1 period, what would it convert to? :D Let's subtract some 360's from it :O

zepdrix (zepdrix):

960 - 2(360) = 240ish? Ok we need to remember our trig identity for this one :) Hmm So we want to find:\[\frac{ 1 }{ \sin 240^o }\] Is sin(240) a special angle? :O Hmmmm it might be!!

zepdrix (zepdrix):

Where you at shew? :D

OpenStudy (anonymous):

i'm typing it in my calculator

OpenStudy (anonymous):

alright i figured it out i got it thanks